OpenStudy (anonymous):

solve (dy/dx)^2-x dy/dx+y=0

OpenStudy (anonymous):

Maybe You could use the equation tool to express yourself more clearly :)

OpenStudy (anonymous):

@BSwan help him

OpenStudy (anonymous):

y''-xy'+y=0 i hate partial equations xD

OpenStudy (anonymous):

power series solution o_o

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

i may be correct or not. i denote \[\frac{ dy }{ dx }=y \prime,~and~\frac{ d^2y }{ dx^2 }=y \prime \prime \] \[\left( y' \right)^2-x y'+y=0\] differentiate again \[2 y'y \prime \prime -xy \prime \prime -y'+y'=0,\] \[\left( 2y \prime-x \right)y \prime \prime=0\] either \[y \prime \prime=0,\frac{ d^2y }{ dx^2 }=0,\] integrating \[\frac{ dy }{ dx }=c _{1}\] again integrating \[y=c _{1}x+c _{2}\] \[2y \prime=x,\frac{ dy }{ dx }=\frac{ 1 }{ 2 }x\] separate the variables and integrating \[y=\frac{ 1 }{ 4 }x^2+c _{3}\]

OpenStudy (anonymous):

hats off for u 

OpenStudy (anonymous):

someone posted this problem yesterday but i did not see anywhere. so i posted and wrote the solution which came in my mind.